OpenStudy (anonymous):

Simplify.[5/(4x^2y)]-[y/(14xz)] Someone please help me

OpenStudy (anonymous):

\[\LARGE \cfrac{5}{4x^2y}-\cfrac{y}{14xz}=\] \[\LARGE \cfrac{14xz \cdot 5 }{4x^2y\cdot 14xz }-\cfrac{y\cdot 4x^2y}{14xz\cdot 4x^2y}=\] \[\LARGE =\cfrac{14xz \cdot 5 - y\cdot 4x^2y}{4x^2y\cdot 14xz }=\] \[\LARGE =\cfrac{x(14z \cdot 5 - y\cdot 4xy)}{x^3y\cdot 56z }=\] \[\LARGE =\cfrac{70z - y^2\cdot 4x}{x^2y\cdot 56z }\] I don't know if this helps... I would call it more: "complicating" than "simplifying" ... lol

OpenStudy (anonymous):

no, this works. thnk u so much, i appreciate u taking the time to reply

OpenStudy (anonymous):

My pleasure :)